The monthly text plan varies inversely as the number of text messages sent. If I send 450 messages for 80 cents a message, what is an inverse variation equation I can use that relates the per message rate and the number of text messages?
Answer by stanbon(75887)

The monthly text plan varies inversely as the number of text messages sent.
p = k/n
Solve for "k" using "If I send 450 messages for 80 cents a message, what is an inverse variation equation?"
----
80 = k/450
k = 36000
---
Equation:
plan = 36000/(number of message)
=====================================
Cheers,
Stan H.